ext Problem 8 The country of Syvania produces and consumes only three goods: Red Bull, pizza, and Tshirts. The quantty produced and price of each good in 2011 and 2012 are given in the following table 2011 2012 Quantity Price Quantity Price 80 540 1020 526 $2 $3 $26 T-Shirts Red Bull (cans) Pizza (alices) 95 540 040 $6 Nominal GOP in 2011 was and nominal GDP in 2012 was sEnter your responses as intepers)
Answer #1

Nominal GDP in 2011 = (80)(26) + (540)(2) + (1020)(3)= $(2080 +1080+3060)= $ 6220

Nominal GDP in 2012 = (96)(26) +(540)(3) +(940)(6)= $(2496+ 1620+ 5640)=$ 9756